OpenStudy (joftheworld):

i need help with exponential function problems please .. ill give medals and fan..

OpenStudy (jdoe0001):

2 ducks 3 pelicans and 11 seagulls

OpenStudy (jdoe0001):

tis easier if you post the specifics... if anyone knows.. .they'd help

OpenStudy (anonymous):

what are the problems

OpenStudy (joftheworld):

@TaylorS11

OpenStudy (anonymous):

it would have to be at least 55 million because it grew 3 million within 10 years

OpenStudy (anonymous):

about 54.5

OpenStudy (joftheworld):

i dont get it.. can you explain more ?

OpenStudy (anonymous):

okay, the population grew 3 million within 10 years, do you get it?

OpenStudy (joftheworld):

no i get that.. but it says its incorrect..

OpenStudy (anonymous):

what are the answer choices

OpenStudy (joftheworld):

theres no answer choices.. i have to type it.. and its quite sensitive..

OpenStudy (joftheworld):

oh okay never mind its not looking for decimals.. its looking for millions so i type 55

OpenStudy (anonymous):

that should be right if you rounded up, was it?

OpenStudy (joftheworld):

well it said i was correct..

OpenStudy (anonymous):

Brilliant!!

OpenStudy (anonymous):

Is that it? You said Problems so if you have any more questions let me know! ;)

OpenStudy (anonymous):

Okay, so from 1995 to 2002 how much did the population grow?

OpenStudy (joftheworld):

7 years?

OpenStudy (anonymous):

yes, and from 2002 to 2018, how many years is that?

OpenStudy (joftheworld):

it growed 4million each 7 years

OpenStudy (joftheworld):

16 years from 2002 to 2018

OpenStudy (anonymous):

So the growth rate is4 million per 7 years

OpenStudy (anonymous):

understand?

OpenStudy (joftheworld):

yeah

OpenStudy (joftheworld):

so they raise 217million by 2018 ?? or am i off ?

OpenStudy (anonymous):

no, it is 4 million for 7 years

OpenStudy (anonymous):

7 goes into 16 twice, so 8 million, but we have a leftover 2 years, so 9 million

OpenStudy (joftheworld):

so 209million ??

OpenStudy (joftheworld):

and since i have to round to the nearest million its 210million ??

OpenStudy (anonymous):

No, 201 million plus 9 million equals

OpenStudy (anonymous):

no, it would be 8 million, but since 14 years is still 2 away from 16, we estimate that it is one million, so 9 million plus the original 201 million

OpenStudy (joftheworld):

next one is this.. but im gonna try it by my self.. just tell me if im making any mistakes..

OpenStudy (anonymous):

okay go ahead

OpenStudy (joftheworld):

so every 6 years they raise 5 million ??

OpenStudy (joftheworld):

but im stuck at adding up the years right.. because 1999 + 6 +6 = 2011.. when its looking for 2012..

OpenStudy (joftheworld):

ill brb i have to eat.. ill be back in a few minutes or so.. please stay here.. i need as much help as possible.. thank you :)

OpenStudy (anonymous):

lol i just ate that why i was gone soory

OpenStudy (joftheworld):

im back lol snack break

OpenStudy (anonymous):

ok you are correct, 5 million every 6 years, for 12 years it would be 10 million, and we would end up in 2011. so we would add 1 million for the remaining year to end up in 2012. Do you follow?

OpenStudy (anonymous):

12 years:2011:10million 13 years:2012:11million

OpenStudy (joftheworld):

so then itll be 187million by 2012

OpenStudy (anonymous):

YES!!! Brilliant!

OpenStudy (anonymous):

So do you understand now?

OpenStudy (joftheworld):

yes thank you.. can you help me with a few other problems?

OpenStudy (anonymous):

of course!

OpenStudy (joftheworld):

this one seems a little difficult..

OpenStudy (anonymous):

Let's use growth rate, \[\frac{ 3300-3000 }{ ?8-0}\]

OpenStudy (anonymous):

\[\frac{ 300 }{ 8}\] then solve: \[300\div8=37.5\]

OpenStudy (joftheworld):

so would it be 38?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

37.5 is the rate of change per year.

OpenStudy (anonymous):

\[37.5\times12=450\]

OpenStudy (anonymous):

So the bacteria would be at 3,450

OpenStudy (anonymous):

Do you understand?

OpenStudy (joftheworld):

yes thanks.. next one is

OpenStudy (anonymous):

Do you remember how I solved the other one?

OpenStudy (joftheworld):

1200 every 7 hours

OpenStudy (anonymous):

No, use rate of change

OpenStudy (anonymous):

\[\frac{ 5200-4000 }{ 7-0 }\]

OpenStudy (anonymous):

\[\frac{ 1200 }{ 7 }\]

OpenStudy (joftheworld):

hmm.. 4000-1200/7 = 400

OpenStudy (joftheworld):

whoops 2,800 sorry

OpenStudy (anonymous):

no 1200/7

OpenStudy (anonymous):

171.4

OpenStudy (anonymous):

So now what do we do?

OpenStudy (joftheworld):

171.4 * 9 ?

OpenStudy (anonymous):

Yes exactly!!

OpenStudy (joftheworld):

1,542.6 so round it to 1,543 ?

OpenStudy (anonymous):

yes you are doing brilliantly

OpenStudy (joftheworld):

it said it was incorrect...

OpenStudy (anonymous):

That s strange try 1,548

OpenStudy (joftheworld):

no.. that came out incorrect aswell..

OpenStudy (anonymous):

IDK what we did wrong

OpenStudy (joftheworld):

ill skip to the next one..

OpenStudy (anonymous):

1800 per two hours or 900 per hour

OpenStudy (joftheworld):

900 per hour and 1800 per 2 hours.

OpenStudy (joftheworld):

so 13,500 for 15 hours

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Okay I have to go so these people will help you! I'm so sorry! But you have been doing brilliantly! @HELP!!!! @superhelp101 @SolomonZelman @undeadknight26 @KlOwNlOvE

OpenStudy (joftheworld):

it says its incorrect D:

OpenStudy (anonymous):

I'm so sorry but those people will help you bc i have to go!

OpenStudy (joftheworld):

@superhelp101 any idea what we did wrong on the last problem ?

OpenStudy (anonymous):

hmm i wouldnt use that formula id use arithmetic sequence because you basically have information for it and its simple

OpenStudy (anonymous):

but first we need to find the growth per hour

OpenStudy (anonymous):

7800-6000=1800/2=900

OpenStudy (joftheworld):

900 per hour. 1800 per 2 hours

OpenStudy (superhelp101):

okay, i think you got this @KlOwNlOvE !! :)

OpenStudy (anonymous):

@superhelp101 would it still work? or does that formula need to be used

OpenStudy (superhelp101):

well i am not sure but i use the formula

OpenStudy (anonymous):

you use that one ill use arithmetic and compare answers?

OpenStudy (anonymous):

a1+(n-1)d a1=starting point n=solving for d=difference

OpenStudy (superhelp101):

okay after that i'll show you what i get and we will se if our answers are right !